Bradley Stoke Community School

Frog Update

We are in the process of transferring Frog over to a new server in order to address the connectivity issues that everyone has been experiencing outside of school.

Starting at 4pm today (Tuesday 21st May) the first stages of data transfer from the old server to the new one will begin.  This will run through the night and although Frog will still be available during this time users will experience a dip in performance.

Between 9-11am tomorrow morning (Wednesday 22nd May) Frog will be unavailable whilst the final data transfer and switch to the new server is complete.

We will post a further update tomorrow once this final process is complete.

Apologies for the inconveniences this is causing